OpenStudy (anonymous):

Locate and classify all the critical points of the function: h(x, y) = 9x^2y − 2x^2 − 4y^2

OpenStudy (anonymous):

The gradient of h is \[\left\{18 x y-4 x,9 x^2-8 y\right\} \] The gradient is zero at \[ p1=( 0, 0) \\ p2=\left( -\frac{4}{9} , \frac{2}{9} \right)\\ p3=\left( \frac{4}{9} , \frac{2}{9} \right)\\ \] Compute the Hessian Determinant at each point and try to classify these points

OpenStudy (anonymous):

and also the second derivative test'

OpenStudy (anonymous):

The hessian Determinant at (ab) is defined as \[h_{\text{xx}}(a,b) h_{\text{yy}}(a,b)-h_{\text{xy}}^2(a,b) \]

OpenStudy (anonymous):

In your particular function h, The hessian Determinant at (a,b) is \[ -324 a^2-144 b+32 \]

OpenStudy (anonymous):

Look how your graph look near (0,0)

OpenStudy (anonymous):

Near the point \[ \left ( \frac{4}{9}, \frac{2}{9} \right) \] The graph looks like the attached graph.